Calculate the discriminant of 9x^2+5x-11=0

This is what I already tried:
5x^2-4(9)(-11)
25-4(9)(-11)
25-396
-371

Answer by Alan3354(69443)

Calculate the discriminant of 9x^2+5x-11=0
------------
a = 9
b = 5
c = -11
-------
Disc = b^2 - 4ac = 25 + 396
Disc = 421
----------
You subtracted the 396. -4*9*-11 = +396